Convert the following decimal fractions to binary with a maximum of six places to the right of the binary point. Use the subtraction method.

(a) 53.8125
Answer #1

in order convert 53.8125 to binary convert 53 and .8125 seperately

if we convert 53 to binary we get (110101)2  

now convert .8125 to binary

.8125*2=1.625-1

.625*2=1.250-1

.250*2=.500-0

.500*2=1.00-1

0.0

binary reperesentation of .8125 is(1101)2

on combing both 53.8125=(110101.1101)2

.